OpenStudy (anonymous):

I need help with solving these systems for the x,y,and z. I know the basic steps but I need help working it out! x - 3y + 4z = -9 2x + y + 3z = 6 x - 4y + 3z = -9 Please do not just give me the answer I need to learn this.

OpenStudy (anonymous):

@ChmE pleaseee

OpenStudy (anonymous):

Go through the same steps as http://openstudy.com/study#/updates/50b7ff75e4b0c789d50ff722

OpenStudy (anonymous):

I'll check your answer

OpenStudy (anonymous):

okay thanks I need a second though

OpenStudy (anonymous):

which method would i use substitution? elimination?

OpenStudy (anonymous):

elimination til you get down to 2 equations. Then I prefer to substitute. But you could eliminate again to solve for one variable.

OpenStudy (anonymous):

I started out by multiplying the first and third equation by 2 is that right

OpenStudy (anonymous):

keep going

OpenStudy (anonymous):

the equations I got to start with are 1.2x-6y+8z=-18 2.2x+y+3z=6 3.2x+8y+6z=18 And then I added ( because my instructor told me to add only) 1. & 2. and got: 4x-5y+11z=-12, but I don't understand I think I have to subtract to get x by its self

OpenStudy (anonymous):

when I subtract I get -7y+5z=-24

OpenStudy (anonymous):

multiply the 1st equation by -1. Then add equations one and two

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

x - 3y + 4z = -9 2x + y + 3z = 6 x - 4y + 3z = -9 2x - 6y + 8z = -18 2x + y + 3z = 6 2x - 8y + 6z = -18 2x - 6y + 8z = -18 -2x + -y + -3z = -6 -7y + 5z = -24 2x - 6y + 8z = -18 -2x + 8y + -6z = 18 2y + 2z = 0 -7y + 5z = -24 2y + 2z = 0 -7y + 5z = -24 y = -z -7(-z) + 5z = -24 7z + 5z = -24 12z = -24 z = -2 y = -z y = -(-2) y = 2 x - 3y + 4z = -9 x - 3(2) + 4(-2) = -9 x - 6 - 8 = -9 x - 14 = -9 x = 5 y = 2 z = -2

OpenStudy (anonymous):

In the 4th step how did you get a positive 18

OpenStudy (anonymous):

oh u multiplied by 1

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

that is from the third equation which was multiplied by -1, correct

OpenStudy (anonymous):

Okay I just worked it out and it matched thanks so much!Again lol,

OpenStudy (anonymous):

yup. You should be able to do all the rest now. I, like your teacher, prefer adding. It is harder to make a stupid mistake with all the negative signs floating around